# Tim ${}^{\dagger}$ 1122 $0$-0002-7305-2667    Markus Ryll 22 $0$-0003-2203-2946 Alex ${}^{\star}$

### v1

@@ Abstract
Visual relationship detection aims to identify objects and their relationships in images. Prior methods approach this task by adding separate relationship modules or decoders to existing object detection architectures. This separation increases complexity and hinders end-to-end training, which limits performance. We propose a simple and highly efficient decoder-free architecture for open-vocabulary visual relationship detection. Our model consists of a Transformer-based image encoder that represents objects as tokens and models their relationships implicitly. To extract relationship information, we introduce an attention mechanism that selects object pairs likely to form a relationship. We provide a single-stage recipe to train this model on a mixture of object and relationship detection data. Our approach achieves state-of-the-art relationship detection performance on Visual Genome and on the large-vocabulary GQA benchmark at real-time inference speeds. We provide analyses of zero-shot performance, ablations, and real-world qualitative examples.
${}^{\dagger}$ork done while at Google DeepMind.
${}^{\star}$dvising project leads in alphabetical order.
@@ 1 Introduction
Our method achieves state-of-the-art visual relationship detection performance on the Visual Genome dataset (29.5% mR@100) and on the large-vocabulary GQA benchmark in open-vocabulary and zero-shot settings. It provides strong results while being significantly simpler than prior approaches to visual relationship detection.
@@ 2 Related Work
Object classes in the natural world follow a long-tailed distribution [12, 62], and this is compounded in VRD due to the combination of subject, object, and predicate in a relationship [27, 49, 48]. A large number of VRD methods specifically aim to address this issue, for example with debiasing losses [52], data augmentation [57], or data resampling [8]. Additionally, while VRD models historically assumed a fixed set of object and predicate classes, recent works add open-vocabulary capabilities [13, 41, 61, 54]. These models leverage pretrained vision-language models such as CLIP [37, 18, 51, 56] to inherit their natural-language understanding. Instead of learning a fixed set of classes, open-vocabulary models predict class embeddings that can be compared to the text embeddings of arbitrary object or predicate descriptions for classification. Here, we employ a pretrained vision-language backbone and combine it with a carefully designed, lightweight relationship detection head to preserve and transfer semantic knowledge from the backbone pretraining to VRD. Without adding further complexity to handle rare classes, this allows our model to achieve strong results on the large-vocabulary QGA benchmark [16] (Section 4.3).
@@ 3.2 Extension to Relationship Prediction
In the architecture described above, each image encoder token represents an object proposal that captures per-object information. Importantly, the encoder consists of self-attention layers which, by design, model all pairwise interactions between these tokens. We therefore hypothesize that information about object relationships can be learned directly in the image encoder. To extract this information in the form of <subject-predicate-object> triplets, we introduce two MLPs that transform vision encoder output tokens into <subject> embeddings $\mathbf{s}_{i}$ and <object> embeddings $\mathbf{o}_{j}$ (Figure 2). Using different MLPs for <subject> and <object> is crucial to break symmetry in subsequent processing.
An embedding representing the relationship between two object proposals (encoder tokens) can then be obtained simply by element-wise addition of their <subject> and <object> embeddings. Obtaining relationship embeddings for all $N^{2}$ <subject-object> pairs, where $N$ is the number of object proposals, would be computationally infeasible. We therefore introduce a Relationship Attention layer that performs hard attention to select the <subject-object> pairs most likely to form a relationship. This layer computes an attention-like score $p_{ij}=\mathbf{s}_{i}\mathbf{o}_{j}^{T}$, where $\mathbf{s}_{i}$ and $\mathbf{o}_{j}$ are the embedding vectors of <subject> $i$ and <object> $j$, and $p_{ij}$ represents the likelihood that a relationship between <subject> $i$ and <object> $j$ exists. By computing this score for all <subject-object> pairs, we obtain an $N{\times}N$ matrix, from which we select the top $k$ pairs. [^2] This selection reduces the number of relationships that need to be processed e.g. from $N^{2}=3600^{2}$ to $k=16386$ (99.9% reduction) for our ViT-L/14 model. For these $k$ pairs, we compute relationship embeddings $\mathbf{r}_{ij}=\textrm{LayerNorm}(\mathbf{s}_{i}+\mathbf{o}_{j})$. To classify the relationship of a pair, its relationship embedding is compared to text embeddings of relationship predicates. We use the embeddings where <subject> and <object> are identical ($\mathbf{r}_{i=j}$) to represent object instances, and predict object classes from them. Boxes are predicted from the corresponding image encoder output tokens (see Figure 2).
@@ 4.1 Experimental Setup
Visual Genome [26], is the largest VRD dataset, labeled with 2.3M triplet relationships across 108K images. However, as Visual Genome includes noisy annotations, the community commonly evaluates VRD on a cleaned version of the dataset [55] where the label space is reduced to the 150 most frequent object classes and the 50 most frequent predicate classes. This dataset is commonly referred to as Visual Genome 150 (VG150).
@@
Unless otherwise noted, we train models on a mixture of VG, VG150, GQA200, HICO, and O365 in the proportions show in Table 1. The B/32 model is trained for 200’000 steps at batch size 256 (further details in Section 0.A.1). Since the datasets we use share similar image sources, there is some overlap between the official training and evaluation splits. To ensure that no evaluation images are used for training, we rigorously filter all training datasets to remove images present in any of the test splits for all datasets we evaluate on, i.e. VG150, GQA, GQA200, and HICO. We use an image similarity filter that also identifies non-exact matches [24]. Table 1 shows the number of images removed from each training split. Note that all prior work performs similar deduplication.
VG150
VG
HICO
GQA200
GQA
Objects365
% in training mixture
$12.5\%$
$12.5\%$
$12.5\%$
$12.5\%$
$0\%$
$50\%$
# removed
$6587$
$17727$
$729$
$15361$
$18801$
$15306$
@@
Prior work presents a diverse range of evaluation metrics often tailored to specific datasets. For non-exhaustively labelled datasets, like Visual Genome and GQA, using precision-based metrics is inherently inconclusive. Thus, the community has adapted a Recall@$K$ metric for these datasets, where $K$ denotes a fixed budget of <subject-predicate-object> triplets on which the recall is computed. However, this metric is biased towards the more frequent <predicate> classes in the dataset [46, 6]. Therefore, recent approaches use mean Recall@$K$, which calculates Recall@$K$ separately for each <predicate> class in the test data and then averages the results.
@@
Further, evaluation can be “graph-constrained”, where only a single prediction can be made per detected object pair, or “graph-unconstrained”, where any number of predictions can be made per pair. We report unconstrained results in the main paper and results with graph constraint in Section 0.A.2.
@@ 4.2 Relationship Detection Performance
We use Visual Genome as our main benchmark for relationship detection because it is widely adopted in the literature. Table 2 shows results for our models and prior work. Our method significantly improves over the prior best method DT2-ACBS [8] by 5.1 points mR@100 (29.5 vs. 24.4). We also note the large difference in performance to UniVRD [61], which builds on the same detection architecture as our method, but adds a Transformer decoder-based relationship module. To disentangle whether the improvements over UniVRD are due to differences in architecture or training data, we also trained our B/32 model on the UniVRD data mixture, and still observe a large improvement (23.9% VG mR@100 for our method vs. 12.1% for UniVRD when trained on the same data; Table 5). This result suggests that our encoder-only architecture is better suited for VRD than prior decoder-based architectures [61, 7, 20, 29].
@@
SG-Transformer [52] + IETrans + Rwt [57] ResNeXt-101 $16.2$ $18.8$
@@
SG-ViT CLIP: ViT-B/32 $20.5$ $24.8$
SG-ViT CLIP: ViT-B/16 $21.4$ $26.6$
SG-ViT CLIP: ViT-L/14 $\bf{23.9}$ $\bf{29.5}$
Table 2: Performance on the Visual Genome test set. Best results are highlighted in bold.
Table 3: Performance on the GQA200 test set. Best results are shown in bold.
@@
SG-ViT (CLIP: ViT-B/32) $21.9$ $26.1$
SG-ViT (CLIP: ViT-B/16) $23.2$ $27.4$
SG-ViT (CLIP: ViT-L/14) $\mathbf{26.7}$ $\mathbf{32.8}$
@@
SG-ViT (CLIP: ViT-B/32) $9.5$ $11.5$
SG-ViT (CLIP: ViT-B/16) $10.1$ $11.7$
SG-ViT (CLIP: ViT-L/14) $\mathbf{13.5}$ $\mathbf{16.6}$

### v2

@@ Abstract
Visual relationship detection aims to identify objects and their relationships in images. Prior methods approach this task by adding separate relationship modules or decoders to existing object detection architectures. This separation increases complexity and hinders end-to-end training, which limits performance. We propose a simple and highly efficient decoder-free architecture for open-vocabulary visual relationship detection. Our model consists of a Transformer-based image encoder that represents objects as tokens and models their relationships implicitly. To extract relationship information, we introduce an attention mechanism that selects object pairs likely to form a relationship. We provide a single-stage recipe to train this model on a mixture of object and relationship detection data. Our approach achieves state-of-the-art relationship detection performance on Visual Genome and on the large-vocabulary GQA benchmark at real-time inference speeds. We provide ablations, real-world qualitative examples, and analyses of zero-shot performance.
Work done while at Google DeepMind.
Advising project leads in alphabetical order.
@@ 1 Introduction
Our method achieves state-of-the-art visual relationship detection performance on the Visual Genome dataset (26.1% mR@100 with graph constraint) and on the large-vocabulary GQA benchmark. It provides strong results while being significantly simpler than prior approaches to visual relationship detection.
@@ 2 Related Work
Object classes in the natural world follow a long-tailed distribution [12, 62], and this is compounded in VRD due to the combination of subject, object, and predicate in a relationship [27, 49, 48]. A large number of VRD methods specifically aim to address this issue, for example with debiasing losses [52], data augmentation [57], or data resampling [8]. Additionally, while VRD models historically assumed a fixed set of object and predicate classes, recent works add open-vocabulary capabilities [13, 41, 61, 54]. These models leverage pretrained vision-language models such as CLIP [37, 18, 51, 56] to inherit their natural-language understanding. Instead of learning a fixed set of classes, open-vocabulary models predict class embeddings that can be compared to the text embeddings of arbitrary object or predicate descriptions for classification. Here, we employ a pretrained vision-language backbone and combine it with a carefully designed, lightweight relationship detection head to preserve and transfer semantic knowledge from the backbone pretraining to VRD. This allows our model to achieve strong results on the large-vocabulary GQA benchmark [16] (Section 4.3).
@@ 3.2 Extension to Relationship Prediction
In the architecture described above, each image encoder token represents an object proposal that captures per-object information. Importantly, the encoder consists of self-attention layers which, by design, model all pairwise interactions between these tokens. We therefore hypothesize that information about object relationships can be learned directly in the image encoder. To extract this information in the form of <subject-predicate-object> triplets, we introduce two MLPs that transform vision encoder output tokens into <subject> embeddings $\mathbf{s}_{i}$ and <object> embeddings $\mathbf{o}_{j}$ (Figure 2). Using different MLPs for <subject> and <object> is crucial to break symmetry in subsequent processing (Section 0.A.1).
An embedding representing the relationship between two object proposals (encoder tokens) can then be obtained simply by element-wise addition of their <subject> and <object> embeddings. Obtaining relationship embeddings for all $N^{2}$ <subject-object> pairs, where $N$ is the number of object proposals, would be computationally infeasible. We therefore introduce a Relationship Attention layer that performs hard attention to select the <subject-object> pairs most likely to form a relationship. This layer computes an attention-like score $p_{ij}=\mathbf{s}_{i}\mathbf{o}_{j}^{T}$, where $\mathbf{s}_{i}$ and $\mathbf{o}_{j}$ are the embedding vectors of <subject> $i$ and <object> $j$, and $p_{ij}$ represents the likelihood that a relationship between <subject> $i$ and <object> $j$ exists. By computing this score for all <subject-object> pairs, we obtain an $N{\times}N$ matrix, from which we select the top $k$ pairs. [^2] This selection reduces the number of relationships that need to be processed, e.g., from $N^{2}=3600^{2}$ to $k=16386$ (99.9% reduction) for our ViT-L/14 model. For these $k$ pairs, we compute relationship embeddings $\mathbf{r}_{ij}=\textrm{LayerNorm}(\mathbf{s}_{i}+\mathbf{o}_{j})$. To classify the relationship of a pair, its relationship embedding is compared to text embeddings of relationship predicates. We use the embeddings where <subject> and <object> are identical ($\mathbf{r}_{i=j}$) to represent object instances, and predict object classes from them. Boxes are predicted from the corresponding image encoder output tokens (see Figure 2).
@@ 4.1 Experimental Setup
Visual Genome [26] is the largest VRD dataset, labeled with 2.3M triplet relationships across 108K images. However, as Visual Genome includes noisy annotations, the community commonly evaluates VRD on a cleaned version of the dataset [55] where the label space is reduced to the 150 most frequent object classes and the 50 most frequent predicate classes. This dataset is commonly referred to as Visual Genome 150 (VG150).
@@
Unless otherwise noted, we train models on a mixture of VG, VG150, GQA200, HICO, and O365 in the proportions show in Table 1. The B/32 model is trained for 200’000 steps at batch size 256 (further details in Section 0.A.1). Since the datasets we use share similar image sources, there is some overlap between the official training and evaluation splits. To ensure that no evaluation images are used for training, we rigorously filter all training datasets to remove images present in any of the test splits for all datasets we evaluate on, i.e. VG150, GQA, GQA200, and HICO. We use an image similarity filter that also identifies non-exact matches [24]. Table 1 shows the number of images removed from each training split. Note that not all prior work performs similar deduplication.
VG150 VG HICO GQA200 GQA Objects365
% in training mixture $12.5\%$ $12.5\%$ $12.5\%$ $12.5\%$ $0\%$ $50\%$
# removed $6587$ $17727$ $729$ $15361$ $18801$ $15306$
@@
Prior work presents a diverse range of evaluation metrics, often tailored to specific datasets. For non-exhaustively labelled datasets, like Visual Genome and GQA, using precision-based metrics is inherently inconclusive. Thus, the community has adapted a Recall@$K$ metric for these datasets, where $K$ denotes a fixed budget of <subject-predicate-object> triplets on which the recall is computed. However, this metric is biased towards the more frequent <predicate> classes in the dataset [46, 6]. Therefore, recent approaches use mean Recall@$K$, which calculates Recall@$K$ separately for each <predicate> class in the test data and then averages the results.
@@
Further, evaluation can be “graph-constrained”, where only a single prediction can be made per detected object pair, or “graph-unconstrained”, where any number of predictions can be made per pair. Existing works evaluating on Visual Genome and GQA datasets partially lack a clear distinction between graph-constrained and unconstrained evaluation. For our models, we report graph-constrained results in the main paper, which reflects the dominant approach used in other studies. Results without graph-constraint are reported in Section 0.A.2.3. For the HICO dataset, we follow the established evaluation procedure [4] that inherently represents a graph-unconstrained evaluation.
@@ 4.2 Relationship Detection Performance
SG-Transformer [52] + IETrans + Rwt [57] ResNeXt-101 $15.0$ $18.1$
@@
SG-ViT CLIP: ViT-B/32 $15.0$ $18.1$
SG-ViT CLIP: ViT-B/16 $15.7$ $19.3$
SG-ViT CLIP: ViT-L/14 $17.8$ $21.8$
SG-ViT + simple predicate rebalancing CLIP: ViT-L/14 $\bf{22.3}$ $\bf{26.1}$
Table 2: Performance on the Visual Genome 150 test set. We use graph-constrained evaluation (one predicate prediction per <subject-object> pair). For the last row (“+ simple text rebalancing”), the model was briefly fine-tuned on VG150 data with rebalanced predicate frequencies (see Sections 4.2 and 0.A.1).
We use Visual Genome as our main benchmark for relationship detection. Table 2 shows results for our models and prior work. Prior studies often achieve reported performance through the combination of a base model with a specialized approach to overcome the skewed data distribution (indicated as base-model + specialization in Table 2). Our models provide strong performance even without special treatment of the skewed predicate distribution. After brief fine-tuning on data in which the frequency of predicate annotations was rebalanced with simple rejection sampling (see Section 0.A.1), our method improves on the prior best method DT2-ACBS [8] by 1.7 points mR@100 (26.1 vs. 24.4). This shows that our already-strong method may be further improved by combining it with more advanced rebalancing methods.
We also note the large difference in performance to UniVRD [61], which builds on the same detection architecture as our method, but adds a Transformer decoder-based relationship module. To disentangle whether the improvements over UniVRD are due to differences in architecture or training data, we also trained our B/32 model on the UniVRD data mixture, and still observe a large improvement (17.4 VG mR@100 for our method vs. 12.1 for UniVRD when trained on the same data; Table 5). This suggests that our encoder-only architecture is better suited for VRD than prior decoder-based architectures [61, 7, 20, 29].
@@
SG-ViT (CLIP: ViT-B/32) $15.7$ $18.6$
SG-ViT (CLIP: ViT-B/16) $17.7$ $20.5$
SG-ViT (CLIP: ViT-L/14) $\mathbf{19.3}$ $\mathbf{22.9}$
Table 3: Performance on the GQA200 test set. Graph-constrained evaluation.
@@
SG-ViT (CLIP: ViT-B/32) $6.2$ $7.4$
SG-ViT (CLIP: ViT-B/16) $6.4$ $7.4$
SG-ViT (CLIP: ViT-L/14) $\mathbf{8.0}$ $\mathbf{9.6}$
Table 4: Performance on the GQA test set. Graph-constrained evaluation.
4.3 Scaling to Large Vocabularies
Improving long-tail performance has been critical to the VRD field since the natural distribution of relationship triplets is highly skewed [27]. A substantial body of literature is devoted to this goal, often proposing complex loss debiasing or data augmentation approaches [27, 52, 57, 8]. It is therefore of interest whether our method works well in this regime without special treatment of rare classes. To evaluate the performance of our method on large object and predicate vocabularies with a long tail of rare classes, we use the Graph Question Answering (GQA) dataset [16]. This dataset builds on VG and expands the number of classes and annotations. A simplified version of GQA with 200 object and 100 predicate classes, called GQA200 [9], is part of our training mixture. Our method surpasses prior results on the GQA200 test split (Table 4), despite lacking the specialized treatment of the data distribution that some prior methods use.
The full GQA dataset has an even larger vocabulary with 1703 object and 311 predicate classes. To our knowledge, no prior methods evaluate scene graph generation on the full GQA dataset. Some prior methods report results on scene graph classification, where ground-truth boxes are provided. As shown in Table 4, our model achieves higher performance on the much harder task of scene graph generation than prior methods on classification. We further assess zero-shot generalization to unseen classes in Section 0.A.2.1.
4.4 Human-Object Interaction
We also evaluate our model on the specialized task of human-object interaction, using the HICO benchmark [4]. The performance of our model is on par with the most comparable prior method [61], but does not show similarly large improvements as for VG or GQA200 (LABEL:tbl:hico). HICO has a much narrower and more specialized vocabulary than VG or GQA (HICO evaluates on 600 pre-specified triplets, whereas VG150 has $1.1\times 10^{6}$ possible triplet combinations). Performance on this specialized task may benefit less from transfer of pretrained representations and may instead be limited by the amount of task-specific training data. This is supported by the fact that the recent state-of-the-art method for HICO pretrains on large amounts of person-focused pseudo-labels [54].
4.5 Ablations
4.5.1 Inference Speed and Number of Predicted Relationships.
The primary hyperparameter of our method is the number of relationships $k$ that are selected by the Relationship Attention layer from all possible <subject-object> combinations for further processing. Figure 3 shows how VRD performance and inference speed depend on $k$.
For training, we use $k=2^{14}$, which results in 71% (B/32) to 81% (L/14) of the speed of a pure object detection model [34] of the same size. Without top-$k$ selection, the model would be several times slower (right end of top plot in Figure 3). Since the vast majority of <subject-object> pairs do not form relationships, VRD performance is essentially unaffected by top-$k$ selection at inference. For inference, we can therefore use $k=2^{11}$ at no loss in accuracy to achieve 90% of the speed of the object detector (for B/32 and L/14). At this $k$, the B/32 model achieves 52.8 FPS at batch size 1 on an NVIDIA V100 GPU and is therefore suitable for real-time applications.
4.5.2 Dataset Mixture.
To assess how training data and architectural improvements contribute to the performance of our model compared to prior work, we train the models on a range of dataset mixtures (Table 5). The full mixture, which we use for all experiments unless otherwise noted, includes VG, VG150, GQA200, HICO and O365. To allow direct comparison to UniVRD [61], we also train on their mixture (VG150, HICO, O365, COCO) and find that our model still performs better on VG150 by a large margin (18.1 mR@100 vs 12.1 for UniVRD for B/32 models). Since UniVRD uses the same image and text encoding backbones as our model, this result suggests that our encoder-only model with Relationship Attention is an advance over decoder-based relationship architectures. We further find that including both pure object detection (O365) and large-vocabulary VRD data (VG) benefit all metrics.
Mixture VG150 GQA200 HICO
VG VG150 GQA200 HICO O365 COCO mR@$100$ mR@$100$ mAP
Our mixture $12.5\%$ $12.5\%$ $12.5\%$ $12.5\%$ $50\%$- $18.1$ $18.6$ $28.9$
UniVRD[61] mix -- $10\%$- $20\%$ $50\%$ $20\%$ $17.4\hskip 2.10002pt\downarrow\hskip 3.5pt4\%$ $\hskip 3.5pt5.1\hskip 2.10002pt\downarrow 73\%$ $17.4\hskip 2.10002pt\downarrow\hskip 3.5pt8\%$
ours w/o O365 $25\%$ $25\%$ $25\%$ $25\%$- -- $15.3\hskip 2.10002pt\downarrow 15\%$ $15.6\hskip 2.10002pt\downarrow 16\%$ $23.8\hskip 2.10002pt\downarrow 18\%$
ours w/o VG -- $25\%$ $12.5\%$ $12.5\%$ $50\%$- $14.2\hskip 2.10002pt\downarrow 22\%$ $15.3\hskip 2.10002pt\downarrow 18\%$ $26.8\hskip 2.10002pt\downarrow\hskip 3.5pt7\%$
Table 5: Dataset Ablation. Performance of SG-ViT (B/32) trained on different data mixtures. VG150 and GQA200 evaluation is graph-constrained, HICO is unconstrained.
4.5.3 Object Detection Performance.
Relation GT SG-ViT (B/32) SG-ViT (L/14)
(a) building near track Refer to caption Refer to caption Refer to caption
(b) light on bus Refer to caption Refer to caption Refer to caption
(c) snow on hill Refer to caption Refer to caption Refer to caption
(d) bottle in door Refer to caption Refer to caption Refer to caption
4.6 Qualitative Results
In Figure 4, we visualize qualitative examples to highlight difficult edge cases. Even where the model output differs from the ground truth, it is often plausibly correct, e.g. where either the extent of an object differs with the ground truth or there are multiple appropriate, but unannotated, relations. For example, both models select the headlight on the bus in Figure 4(b) for “light on bus” as opposed to the sun reflected off the windshield. In Figure 4(c) both models focus more on the <predicate> “on” in “snow on hill” than the distinction between a mountain and a hill. Figure 4(d) shows a challenging scene where the B/32 model selects the bottle inside the open fridge while the L/14 correctly selects a water bottle in the door for “bottle in door” while also capturing the extent of the door instance.
5 Limitations
While our model performs strongly on large-vocabulary VRD, its performance on specialized human-object interaction detection is only on par with prior models. We discuss this limitation in Section 4.4. Further error modes are explored qualitatively in Section 4.6 and Section 0.A.3. A challenge for open-vocabulary relationship detection models as a whole is zero-shot generalization to unseen objects and predicates. While our model improves over prior approaches (Tables 4 and 4), there is still a large gap when it comes to entirely unseen classes (Section 0.A.2.1). Future research should focus on closing this gap.
6 Conclusion
We present an architecture for open-vocabulary visual relationship detection. By combining an encoder-only design with a novel Relationship Attention layer for efficient selection of high-confidence relationships, our architecture leverages VLM pretraining and multi-dataset VRD training. It achieves strong performance on standard and large-vocabulary VRD benchmarks while maintaining pure object detection performance and adding little extra computational cost. Due to its simplicity and strong performance, we believe that our method will be a useful basis for further research on visual relationship detection.
Appendix 0.A Appendix
0.A.1 Model Details and Hyperparameters
0.A.1.1 Relationship Attention Architecture.
Figure 6: Computing <subject> and <object> embeddings with separate MLPs is necessary for good performance.
To obtain query, key, and value embeddings from the image embeddings for the Relationship Attention layer, we use 3-layer MLPs with no change in feature dimensionality, GeLU hidden activations [14], and a skip connection from the input to the output embedding. LayerNorm [2] is applied to the final output in the MLPs. To obtain the relationship embedding, <subject> and <object> embeddings are summed, normalized by LayerNorm, and processed by another 2-layer MLP (not shown in Figure 2). We found model performance to be robust to the details of the Relationship Attention layer, e.g. the hyperparameters of the MLPs, and it may be possible to simplify the design further. However, as noted in the main paper, <subject> and <object> embeddings must be computed with different projections to model the asymmetry between <subject> and <object> in the relationship. If the same projection were used to compute a single embedding to represent both <subject> and <object>, the model could not distinguish between e.g. "person riding horse" and "horse riding person". Figure 6 confirms experimentally that sharing the MLP for subjects and objects performs poorly. With a shared MLP, the model struggles to learn and reaches a low final score.
0.A.1.2 Relationship Selection.
0.A.1.3 Data Augmentation.
[…]
0.A.1.4 Data Rebalancing.
For most of our experiments, we use the training data as-is without special treatment of the skewed object or predicate class distributions. Only for the model on the last line in Table 2, we perform simple rebalancing as follows, to show the potential of combining our method with orthogonal approaches focusing on data rebalancing: We first count the number of occurrences of each predicate in the training set to obtain its frequency. Then, during training, we randomly drop relationship annotations with a probability equal to its frequency (capped at 0.95 for the most frequent predicates).
0.A.1.5 Training Details.
0.A.1.6 Speed Benchmarking.
0.A.2 Additional Experimental Results
0.A.2.1 Zero-shot GQA.
To assess zero-shot generalization to unseen classes, we report the performance on the least-frequent 1503 object and 211 predicate classes in GQA, i.e. those not included in GQA200 and therefore unseen during training (Table 6). Given the large vocabulary and the difficulty of zero-shot predicate classification, we evaluate the model without graph-constraint, allowing four predicate predictions per <subject-object> pair. Although the performance of our model in this scenario is nontrivial, it is significantly lower than on seen classes (Table 4). We therefore suggest using the least-frequent GQA annotations in this manner as a challenging benchmark for future work on zero-shot VRD.
0.A.2.2 Recall@K.
0.A.2.3 Performance without Graph-Constraint.
All results in the main paper for VG150 and GQA are computed with graph constraint. Table 8 provides these results without graph constraint.
Model mR@$50$ mR@$100$